OpenStudy (anonymous):

Simplify this trig. expression:

OpenStudy (anonymous):

\[\sec x \sqrt{\frac{ (1- \sin^2 y)(\sin^2x) }{ (1+\cos^2y)(\tan^2x) }}\]

OpenStudy (anonymous):

\[\sec x \sqrt\frac{(1-\sin^2 y)(\sin^2 x)}{(1+\cos^2 y)(\tan^2 x)}=\frac{1}{\cos x}\sqrt\frac{(\cos^2 y)(\sin^2 x)}{(1+\cos^2 y)(\tan^2 x)}\]\[=\frac{1}{\cos x}\frac{(\cos y)(\sin x)}{\sqrt{1+\cos^2 y}(\tan x)}\][Can you do the cancellation for the rest?]

OpenStudy (anonymous):

Do you mind showing one more step?

OpenStudy (anonymous):